OpenStudy (anonymous):

to i simplify large fractions like 181/990 please??

OpenStudy (anonymous):

a quick way to simplify large fractions??

OpenStudy (amistre64):

there are no "quick" ways ...

OpenStudy (amistre64):

there are trial and error ways for numbers like this

OpenStudy (anonymous):

oh ok...thank you

OpenStudy (amistre64):

if they are both even, /2 after that try the odds /3 /5 /7 /9 ...

OpenStudy (amistre64):

we know 5s end in 5 or 0 so its out

OpenStudy (amistre64):

181/3 = 60 R 1 so 3s are out

OpenStudy (amistre64):

181/4 = 45 R 1

OpenStudy (anonymous):

thanks alot...i probably sound really dumb but i get confused easily :)

OpenStudy (amistre64):

181/7 = 25 R 6 'sok :) if anything it could very well be simplified already

OpenStudy (anonymous):

true say...

OpenStudy (amistre64):

its as low as it can go http://www.wolframalpha.com/input/?i=181%2F990